OpenStudy (anonymous):

Q = Q0*e^(-t/r).. solve for t

OpenStudy (anonymous):

\[Q=Q_0e^{-\frac{t}{r}}\] \[\frac{Q}{Q_0}=e^{-\frac{t}{r}}\] \[\ln(\frac{Q}{Q_0})=-\frac{t}{r}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw